Find the equation of the line.

Through (3,−5)​; perpendicular to y=−10.


1
Expert's answer
2022-02-22T13:48:35-0500

Explanations & Calculations


  • Equation of a line looks like "\small y = mx+c"
  • The line "\small y = -10" is parallel to "\small x- axis" and since our line should be perpendicular to this, it should be parallel to "\small y - axis" hence should have the form "x = p" .
  • Since it lies on the given point, the line turns out to be "\small x = 3" .
  • This line lies on points "\small -\infty \leftarrow (3,3),\,(3,4),\,(3,5),\,(3,6),\,(3,7)\to+\infty"